6. Write a set of mesh current equations in matrix form for the following circuit. ری ای بارد )10 داده دیا ۔ 247 )
Answer #1

doom Céscuit, Vaxt zo le-ti] OAہا۔ Čg = 40 4x + es (2-40% x + Ez =0 33 - 40x4 pla-] + c3 =0 a-acém-<,) + 3 =0 Ez tég-2cm = +8et és a Sopea Meos cacuit, KVL to loop-2, 3 Catto Cz+ Es Leg- (a) CG-) 30 El the ]$[ ***87- 38 tlo siz-Eu = -2145 1 312 + lo